vgalleg5:

In the circle, AB = 42, BC = 18, and CD = 4.

dude:

We need to know where the lines are within the circle, could you upload an image?

Vocaloid:



Vocaloid:

applying this logic to your problem we have: CD(x + CD) = BC(BC + AB) can you try plugging in the appropriate values and solving for x? I am willing to help out if you get stuck ~

Vocaloid:

4(4+x) = (18)(18+42) x = ?

Vocaloid:

hm. distributing and multiplying gives us 16 + 4x = 1080 x = ?

Vocaloid:

good so 266 is your solution